OpenStudy (ashking1):

Which equation represents the line passing through the points (1, 4) and (-2, 13)? x + 3y = 7 x - 3y = 7 3x - y = 7 3x + y = 7

OpenStudy (lgbasallote):

use the two-point formula \[\Large y - y_1 = \frac{y_2 - y_1}{x_2 - x_1} (x - x_1)\] Now substitute the values... \[\Large y - 4 = \frac{13 - 4}{-2 - 1} (x - 1)\] \[\Large y - 4 = \frac{9}{-3} (x-1)\] \[\Large y - 4 = -3(x-1)\] do you know what to do now?